The Ksp for Cr (OH) 3 is 1.2*10^-15 what is its solubility in pure water

Grade:12

1 Answers

Ankur Saha
7 years ago
Let it`s solubility in water is =sCr (OH) 3=Cr(3+)+3OH(-)Ksp =3s"2=1.2*10"(-15)Thereafter s=2*10"(-8)